# Let Network Decide What to Learn: Symbolic Music Understanding Model Based on Large-scale Adversarial Pre-training

### v2

@@ Abstract
Index Terms—  Music Information Retrieval (MIR), Symbolic Music Understanding (SMU), Adversarial Learning, Bidirectional Encoder Representations from Transformers (BERT)
1 Introduction
Music Information Retrieval (MIR) plays a crucial role in various applications, including recommendation systems in music apps and AI agents for music creation. With advancements in computer music, symbolic music—representing music through a structured sequence of notes—has gained significant attention, as most contemporary music is initially created and recorded using symbolic formats like MIDI [1]. Symbolic Music Understanding (SMU) has emerged as a key research direction within MIR, aiming to assist both musicians and enthusiasts in learning, teaching, and creating music.
Given the similarities between symbolic music and natural language, language models have been extensively utilized in SMU. For instance, the Bidirectional Encoder Representations from Transformers (BERT) [2] model has demonstrated promising performance in SMU [3, 4]. A critical factor contributing to the success of current language models, particularly Large Language Models (LLMs), is their ability to leverage vast amounts of unlabeled data during pre-training to learn fundamental data structures and relationships. This pre-training mechanism has proven effective in domains with limited data, such as music [5] and signals [6], thereby enhancing model performance in downstream tasks [7].
Among the prevalent pre-training methods for language models is the Mask Language Model (MLM) [2], particularly in encoder-only architectures. This method randomly masks certain tokens and trains the model to recover them. However, since some tokens cannot be inferred from the given context, training the model on such tasks can introduce bias issues, such as discrimination based on gender, age, or race in Natural Language Processing (NLP) [8]. Research has indicated that these biases can significantly impair model performance in downstream tasks, including classification and generation [9]. Consequently, even in the absence of ethical discrimination issues in music, bias can adversely affect model performance in SMU.
@@ 1 Introduction
However, most current methods addressing bias are limited to the field of NLP and are challenging to transfer to other domains. Many of these methods tackle specific bias issues, such as sexism [10] and regional bias [11], but do not apply similarly in other areas. For instance, data augmentation [12] is a promising approach in NLP. However, due to the less pronounced bias issues in domains like music and signal processing, it remains unclear how to effectively clean, generate, or modify data without sufficient domain-specific knowledge.
To address the bias problem in SMU, we propose the Adversarial-MidiBERT: (1) We design an adversarial mechanism to avoid masking context-free tokens in order to mitigate bias. (2) We introduce a mask fine-tuning method that incorporates random [MASK] tokens during fine-tuning to bridge the data gap between pre-training and fine-tuning, thereby enhancing model performance. (3) Experimental results demonstrate that our method achieves excellent performance across four music understanding tasks, including composer classification, emotion classification, velocity prediction, and melody extraction.
2 Proposed Method
Fig. 1: Network Architecture
2.1 Overview
@@
Before delving deeper into the training mechanism, we first introduce our network and data structure. To embed MIDI music information, we employ Octuple [3] to represent the symbolic music structure. It transforms each MIDI file into a sequence of tokens, each possessing eight attributes: time signature (TS), tempo (BPM), bar position (BAR), relative position within each bar (POS), instrument, pitch, duration, and velocity. To input these tokens into the network, we utilize eight embedding layers to embed them individually, and then concatenate the embedding results. This process yields an embedding sequence suitable for input to the network, as shown in Fig. 1. In this configuration, a BERT [2] (the encoder component of the Transformer [13]) serves as the backbone, complemented by an embedding layer for music information and multiple output heads for different tasks. Specifically, the recoverer head addresses the MLM task by generating classification probabilities for each attribute at each position. The masker head generates the mask probability for each token. Additionally, the sequence-level and token-level classifiers are employed in downstream classification tasks, producing classification probabilities for the entire music sequence and each token, respectively. Due to page limitations, please refer to [5, 3] for more details regarding the model structure and Octuple information.
Fig. 2: Pre-train Process: The blue tokens represent the frozen tokens, which cannot be selected as [MASK] tokens.
2.2 Adversarial Pre-training
@@
$w_{j}=\frac{\frac{1}{a_{j}}}{\sum_{i=1}^{8}\frac{1}{a_{i}}}\ ,$ (2)
@@
$L_{masker}=\sum_{i\in I_{0}}\text{BCE}(p_{i},0)+\sum_{i\in I_{1}}\text{BCE}(p_{i},1)\ ,$ (3)
@@
2.3 MASK Fine-tuning
@@
3 Experiment
3.1 Experiment Setup
Table 1: Model Configurations
Configuration Our Setting Input Length 1024 Network Layers 12 Hidden Size 768 Inner Linear Size 3072 Attn. Heads 12 Dropout Rate 0.1 Optimizer AdamW Learning Rate $10^{-4}$ (pre-train), $10^{-5}$ (fine-tune) Batch Size 8 Parameters Mentioned in (15,30,30,10,15) Section 2 $(p,q,a,b,k)$ Total Number of Parameters 115 Million
Table 2: Dataset Decription
Dataset Pieces Task Task Level Class Number Used in Pre-training ASAP [14] 1068 Y Pop1K7 [15] 1747 Y Pianist8 [16] 865 Composer Classification Sequence Level 8 Y EMOPIA [17] 1078 Emotion Recognition Sequence Level 4 Y POP909 [18] 909 Melody Extraction Token Level 3 Y GiantMIDI [19] 10855 Velocity Prediction Token Level 6 N
Table 3: Model Performance in Different Tasks: The bold and underlined value indicates the best and second best result within each task. (Note that the training times provided are for reference only, as the server is shared with other users.)
Model Pre-train Sequence-Level Classification Token-Level Classification Accuracy Epochs Time Composer Emotion Velocity Melody MidiBERT [4] 79.60% 500 6.44d 79.07% 67.59% 44.88% 92.53% MusicBERT-QM [20] 80.57% 500 9.47d 83.72% 69.52% 46.71% 92.64% MusicBERT [3] 76.01% 500 10.06d 86.05% 71.06% 38.79% 92.47% PianoBART [5] 96.67% 268 3.19d 88.37% 73.15% 49.37% 92.62% Adversarial-MidiBERT (ours) 81.47% 436 9.82d 97.92% 79.46% 45.58% 92.68% Adversarial-MidiBERT (fine-tune w/o mask) 81.47% 436 9.82d 65.98% 70.53% 45.30% 92.55% Adversarial-MidiBERT (pre-train w/o adversary) 83.51% 500 5.93d 88.91% 74.07% 45.44% 92.64% Adversarial-MidiBERT (w/o pre-train) 79.76% 68.75% 38.70% 87.98%
@@ 3.1 Experiment Setup
    Emotion Recognition: The music emotions in EMOPIA [17] are divided into four types: HVHA, HVLA, LVHA, and LVLA. This task requires the model to classify each song into one of these types.
@@
    Velocity Prediction: Since many MIDI files do not include velocity information, it is important to predict the velocity. We divide velocity into six types and train the model to predict it. To avoid information leakage, we use GiantMIDI [19], which does not participate in pre-training and has its velocity information masked. Since our device could not support training using the full dataset, we select only the first 1000 pieces for the experiment.
We split each dataset into 80% training set, 10% validation set, and 10% testing set. We employ the same early stopping strategy as previous works [5, 4], where the training would stop if the model’s accuracy does not increase on the validation set for 30 consecutive epochs. We also set the maximum training epochs to 500.
3.2 Experiment Result
In this section, we compare our method with other SMU models, including MidiBERT [4], MusicBERT [3], MusicBERT-QM [20], and PianoBART [5]. The main differences between these methods lie in their pre-training approaches. For fairness, these BERT-based models use the same backbone as our method. The experimental results are shown in Table 3. It can be seen that our method outperforms the previous BERT-based methods in most tasks, but loses to PianoBART in pre-training and velocity prediction. This may be influenced by the model structure, as the auto-regressive mechanism of the decoder structure in Bidirectional and Auto-Regressive Transformers (BART) [21] makes it more suitable for token-level tasks. However, it can be noticed that our model has an extremely significant increase in performance on sequence-level tasks. What’s more, we also conducted an series of ablation study to illustrate our method’s performance without pre-training, pre-training without the adversarial-learning mechanism and without masking during fine-tuning. The results show that the model’s performance decreases to varying degrees in these cases, demonstrating the effectiveness of our proposed mechanisms.
4 Conclusion
In this paper, we present Adversarial-MidiBERT for SMU, addressing the bias problem of pre-trained models in MIR through an adversarial pre-training method. Additionally, we introduce a mask fine-tuning approach that significantly enhances the model’s accuracy on downstream tasks. Our method achieves remarkable performance on four SMU tasks, particularly on sequence-level tasks.

### v3

@@ Abstract.
Music Information Retrieval (MIR), Symbolic Music Understanding (SMU), Adversarial Learning, Bidirectional Encoder Representations from Transformers (BERT)
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ACM Conference; June 30 - July 3, 2025; Chicago, USAisbn: 978-1-4503-XXXX-X/18/06journalyear: 2025copyright: acmlicensedconference: Proceedings of the 2025 International Conference on Multimedia Retrieval; June 30-July 3, 2025; Chicago, IL, USAbooktitle: Proceedings of the 2025 International Conference on Multimedia Retrieval (ICMR ’25), June 30-July 3, 2025, Chicago, IL, USAdoi: 10.1145/3731715.3733483isbn: 979-8-4007-1877-9/2025/06ccs: Applied computing Sound and music computingccs: Computing methodologies Artificial intelligence
1. Introduction
Music Information Retrieval (MIR) plays a crucial role in various applications, including recommendation systems in music apps and AI agents for music creation. With advancements in computer music, symbolic music—representing music through a structured sequence of notes—has gained significant attention, as most contemporary music is initially created and recorded using symbolic formats like MIDI (Rothstein, 1995). Symbolic Music Understanding (SMU) has emerged as a key research direction within MIR, aiming to assist both musicians and enthusiasts in learning, teaching, and creating music.
Given the similarities between symbolic music and natural language, language models have been extensively utilized in SMU. For instance, the Bidirectional Encoder Representations from Transformers (BERT) (Devlin et al., 2018) model has demonstrated promising performance in SMU (Zeng et al., 2021; Chou et al., 2021). A critical factor contributing to the success of current language models, particularly Large Language Models (LLMs), is their ability to leverage vast amounts of unlabeled data during pre-training to learn fundamental data structures and relationships. This pre-training mechanism has proven effective in domains with limited data, such as music (Liang et al., 2024) and signals (Zhao et al., 2024), thereby enhancing model performance in downstream tasks (Wang et al., 2020b).
Among the prevalent pre-training methods for language models is the Mask Language Model (MLM) (Devlin et al., 2018), particularly in encoder-only architectures. This method randomly masks certain tokens and trains the model to recover them. However, since some tokens cannot be inferred from the given context, training the model on such tasks can introduce bias issues, such as discrimination based on gender, age, or race in Natural Language Processing (NLP) (Garrido-Muñoz et al., 2021). Research has indicated that these biases can significantly impair model performance in downstream tasks, including classification and generation (Gallegos et al., 2024). Consequently, even in the absence of ethical discrimination issues in music, bias can adversely affect model performance in SMU.
@@ 1. Introduction
However, most current methods addressing bias are limited to the field of NLP and are challenging to transfer to other domains. Many of these methods tackle specific bias issues, such as sexism (Bordia and Bowman, 2019) and regional bias (Li et al., 2022), but do not apply similarly in other areas. For instance, data augmentation (Lu et al., 2020) is a promising approach in NLP. However, due to the less pronounced bias issues in domains like music and signal processing, it remains unclear how to effectively clean, generate, or modify data without sufficient domain-specific knowledge. To address the bias problem in SMU, we propose Adversarial-MidiBERT:
  • $\bullet$
    We introduce an adversarial pre-training approach that uses a masker network to selectively mask context-dependent tokens during MLM, avoiding context-free tokens. This reduces bias from dataset-specific patterns, enabling the model to learn robust musical structures like modes and riffs, improving generalization across SMU tasks.
  • $\bullet$
    We propose a mask fine-tuning method that applies random [MASK] tokens during fine-tuning to align pre-training and fine-tuning phases. This enhances model robustness and performance on downstream tasks by maintaining contextual inference capabilities, especially with limited music data.
  • $\bullet$
    Experimental results demonstrate that our method achieves excellent performance across four music understanding tasks, including composer classification, emotion classification, velocity prediction, and melody extraction. Ablation studies confirm the effectiveness of our adversarial and fine-tuning mechanisms, highlighting the model’s versatility for music analysis.
1.1. Overview
@@
Before delving deeper into the training mechanism, we first introduce our network and data structure. To embed MIDI music information, we employ Octuple (Zeng et al., 2021) to represent the symbolic music structure. It transforms each MIDI file into a sequence of tokens, each representing one music note and possessing eight attributes: time signature (TS), tempo (BPM), bar position (BAR), relative position within each bar (POS), instrument, pitch, duration, and velocity. To input these tokens into the network, we utilize eight embedding layers to embed them individually, and then concatenate the embedding results. This process yields an embedding sequence suitable for input to the network, as shown in Fig. 1. In this configuration, a BERT (Devlin et al., 2018) (the encoder component of the Transformer (Vaswani et al., 2017)) serves as the backbone, complemented by an embedding layer for music information and multiple output heads for different tasks. Specifically, the recoverer head addresses the MLM task by generating classification probabilities for each attribute at each position. The masker head generates the mask probability for each token. Additionally, the sequence-level and token-level classifiers are employed in downstream classification tasks, producing classification probabilities for the entire music sequence and each token, respectively. Due to page limitations, please refer to (Liang et al., 2024; Zeng et al., 2021) for more details regarding the model structure and Octuple information.
2. Proposed Method
Figure 1. Network Architecture
Figure 2. Pre-train Process: The blue tokens represent the frozen tokens, which cannot be selected as [MASK] tokens.
2.1. Adversarial Pre-training
@@
(2) $w_{j}=\frac{\frac{1}{a_{j}}}{\sum_{i=1}^{8}\frac{1}{a_{i}}}\ ,$
@@
(3) $L_{masker}=\sum_{i\in I_{0}}\text{BCE}(p_{i},0)+\sum_{i\in I_{1}}\text{BCE}(p_{i},1)\ ,$
@@
2.2. MASK Fine-tuning
@@
3. Experiment
3.1. Experiment Setup
Table 1. Model Configurations
Configuration Our Setting
Input Length 1024
Network Layers 12
Hidden Size 768
Inner Linear Size 3072
Attn. Heads 12
Dropout Rate 0.1
Optimizer AdamW
Learning Rate $10^{-4}$ (pre-train)
$10^{-5}$ (fine-tune)
Batch Size 8
Parameters Mentioned in (15,30,30,10,15)
Section 2 $(p,q,a,b,k)$
Total Number of Parameters 115 Million
Table 2. Dataset Decription
Dataset Pieces Task Task Level Class Number Used in Pre-training
ASAP (Foscarin et al., 2020) 1068 $\checkmark$
Pop1K7 (Hsiao et al., 2021) 1747 $\checkmark$
Pianist8 (joann8512, 2021) 865 Composer Classification Sequence Level 8 $\checkmark$
EMOPIA (Hung et al., 2021) 1078 Emotion Recognition Sequence Level 4 $\checkmark$
POP909 (Wang et al., 2020a) 909 Melody Extraction Token Level 3 $\checkmark$
GiantMIDI (Kong et al., 2022) 10855 Velocity Prediction Token Level 6 $\times$
Table 3. Model Performance in Different Tasks: The bold and underlined value indicates the best and second best result within each task. (Note that the training times provided are for reference only, as the server is shared with other users.)
Model Pre-train Sequence-Level Classification Token-Level Classification Accuracy Epochs Time Composer Emotion Velocity Melody MidiBERT (Chou et al., 2021) 79.60% 500 6.44d 79.07% 67.59% 44.88% 92.53% MusicBERT-QM (Shen et al., 2023) 80.57% 500 9.47d 83.72% 69.52% 46.71% 92.64% MusicBERT (Zeng et al., 2021) 76.01% 500 10.06d 86.05% 71.06% 38.79% 92.47% PianoBART (Liang et al., 2024) 96.67% 268 3.19d 88.37% 73.15% 49.37% 92.62% Adversarial-MidiBERT (ours) 81.47% 436 9.82d 97.92% 79.46% 45.58% 92.68% Adversarial-MidiBERT (fine-tune w/o mask) 81.47% 436 9.82d 65.98% 70.53% 45.30% 92.55% Adversarial-MidiBERT (pre-train w/o adversary) 83.51% 500 5.93d 88.91% 74.07% 45.44% 92.64% Adversarial-MidiBERT (w/o pre-train) 79.76% 68.75% 38.70% 87.98%
@@ 3.1. Experiment Setup
    Emotion Recognition: The music emotions in EMOPIA (Hung et al., 2021) are divided into four types: HVHA, HVLA, LVHA, and LVLA. This task requires the model to classify each song into one of these types.
@@
    Velocity Prediction: Since many MIDI files do not include velocity information, it is important to predict the velocity. We divide velocity into six types and train the model to predict it. To avoid information leakage, we use GiantMIDI (Kong et al., 2022), which does not participate in pre-training and has its velocity information masked. Since our device could not support training using the full dataset, we select only the first 1000 pieces for the experiment.
We split each dataset into 80% training set, 10% validation set, and 10% testing set. We employ the same early stopping strategy as (Liang et al., 2024; Chou et al., 2021), where the training would stop if the model’s accuracy does not increase on the validation set for 30 consecutive epochs. We also set the maximum training epochs to 500.
3.2. Experiment Result
In this section, we compare our method with other SMU models, including MidiBERT (Chou et al., 2021), MusicBERT (Zeng et al., 2021), MusicBERT-QM (Shen et al., 2023), and PianoBART (Liang et al., 2024). The main differences between these methods lie in their pre-training approaches. For fairness, these BERT-based models use the same backbone as our method. The experimental results are shown in Table 3. It can be seen that our method outperforms the previous BERT-based methods in most tasks, but loses to PianoBART in pre-training and velocity prediction. This may be influenced by the model structure, as the auto-regressive mechanism of the decoder structure in Bidirectional and Auto-Regressive Transformers (BART) (Lewis et al., 2019) makes it more suitable for token-level tasks. However, it can be noticed that our model has an extremely significant increase in performance on sequence-level tasks. What’s more, we also conducted an series of ablation study to illustrate our method’s performance without pre-training, pre-training without the adversarial-learning mechanism and without masking during fine-tuning. The results show that the model’s performance decreases to varying degrees in these cases, demonstrating the effectiveness of our proposed mechanisms.
4. Conclusion
In this paper, we present Adversarial-MidiBERT for SMU, addressing the bias problem of pre-trained models in MIR through an adversarial pre-training method. Additionally, we introduce a mask fine-tuning approach that significantly enhances the model’s accuracy on downstream tasks. Our method achieves remarkable performance on four SMU tasks, particularly on sequence-level tasks. In the future, we aim to explore the application of our method to music generation and NLP tasks. Furthermore, additional experiments can be conducted to further elucidate the principles of the proposed adversarial mechanism.